undeadknight26 (undeadknight26):

lost...please help? Let f(x) = x - 2 and g(x) = x^2 - 7x - 9. Find f(g(-1)).

undeadknight26 (undeadknight26):

f(g(-1)) g(-1) = x^2 - 7(-1) - 9 g(-1) = 1 - 7 - 9 g(-1) = -6 - 9 g(-1) = -15 But now i cannot make: f(g(-1)) i get: -15(-15) - 2

undeadknight26 (undeadknight26):

@satellite73 @waterineyes

undeadknight26 (undeadknight26):

The answer for this was -3.